OpenStudy (samara8954):

Sandy works at a clothing store. She makes $7 per hour plus earns 10% commission on her sales. She worked 72 hours over the last two weeks and had a total of $1,679 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? $672 $168 $504 $218

OpenStudy (samara8954):

@AlexandervonHumboldt2

OpenStudy (anonymous):

7(number of hours worked)+\[(7(x))+10(\frac{ 1679 }{ 100 })= money earned\]

OpenStudy (samara8954):

@AlexandervonHumboldt2

OpenStudy (samara8954):

help me @AlexandervonHumboldt2

OpenStudy (anonymous):

\[7(72)+10(16.79)\]\[= 504+167.9\]\[= 671.9\]\[\approx $672\]

OpenStudy (alexandervonhumboldt2):

672 $ @Sadworld i right why you tag me when he helps you? that look as you ignore him

OpenStudy (anonymous):

Oh my bad, I have a small screen so when the Equation tags comes up, I turn blind.

OpenStudy (anonymous):

Sorry, got to double check before answering. Won't happen again.